# Collocation methods for nonlinear differential equations on low-rank manifolds

### v2

@@ 2.1 Orthogonalization of tensor trains
Orthogonal TT representations are fundamental for executing many operations in the TT format. We will use them in this paper to obtain well-conditioned projectors onto the tangent spaces of TT manifolds. Hereafter, we recall an algorithm for orthogonalizing TTs by recursively applying QR-decomposition to TT-core unfoldings. Begin by taking a QR-decomposition of the left unfolding of $C_{1}$
@@
$Y=U_{\leq k}\bm{S}_{k}V_{>k},$ (6)
and utilizing the unfolding matrices of partial products (5) we also have a decomposition of the $k$th unfolding
@@ 3 Projections onto the TT tangent space
is a smooth embedded submanifold of $\mathbb{R}^{n_{1}\times\cdots\times n_{d}}$ [27]. Hence for any tensor $Y\in\mathcal{M}_{\bm{r}}$ we can define the tangent space $T_{Y}\mathcal{M}_{\bm{r}}$ as the collection of (equivalence classes of) velocities of smooth curves on $\mathcal{M}_{\bm{r}}$ passing through $Y$
$T_{Y}\mathcal{M}_{\bm{r}}=\left\{\gamma^{\prime}(s)|_{s=0}\mid\gamma(s)\in C^{1}((-\delta,\delta),\mathcal{M}_{\bm{r}}),\ \gamma(0)=Y\right\}.$ (9)
The tangent space is a vector subspace of $\mathbb{R}^{n_{1}\times\cdots\times n_{d}}$ that linearizes the manifold $\mathcal{M}_{\bm{r}}$ around $Y$. Given a TT representation (3) of $Y$, any element of the tangent space has the form
@@ 3.1 The orthogonal tangent space projector
onto the range and co-range of $\bm{Y}^{\langle k\rangle}$ (see (7)). These projectors act on the matrix space $\mathbb{R}^{(n_{1}\cdots n_{k})\times(n_{k+1}\cdots n_{d})}$. To construct the orthogonal tangent space projector it is convenient to define projectors corresponding to (11) that act on the tensor space $\mathbb{R}^{n_{1}\times\cdots\times n_{d}}$ by
@@ 3.2 Oblique tangent space projectors
Figure 2: A sketch of the low-rank tensor manifold $\mathcal{M}_{\bm{r}}$ and its tangent space $T_{Y}\mathcal{M}_{\bm{r}}$ at the point $Y\in\mathcal{M}_{\bm{r}}$. Also depicted is a tensor $Z\in\mathbb{R}^{n_{1}\times\cdots n_{d}}$ and its orthogonal projection $\widehat{P}_{Y}Z$ and oblique projection $P_{Y}Z$ onto the tangent space $T_{Y}\mathcal{M}_{\bm{r}}$. The orthogonal projection is the best approximation of $Z$ on the tangent space with respect to the Frobenius norm but is impractical compute when $G$ lacks rank structure. The oblique projection is a quasi-optimal approximation of $Z$ on the tangent space that is efficient to compute for any $G$ that can be evaluated entry-wise.
@@
with $P_{\leq 0}=1$, defined for any tensor $Z\in\mathbb{R}^{n_{1}\times\cdots n_{d}}$, is an oblique projector onto the tangent space $T_{Y}\mathcal{M}_{\bm{r}}$.
@@
which we write in the TT format
@@
We remark that the orthogonality of the bases $\bm{U}_{\leq k}$ and $\bm{V}_{>k}$ is not necessary for the construction of oblique tangent space projectors (16). However, imposing such orthogonality is advantageous for the numerical stability and accuracy of the projectors. Next we introduce a special class of oblique tangent space projectors with a cross interpolation property.
@@ 3.3 Interpolatory tangent space projectors
Let us consider oblique tangent space projectors (16) with $\bm{X}_{\leq k}$ and $\bm{X}_{>k}$ defined as $r_{k}$ columns of the identity matrix with appropriate dimension
@@
We obtain interpolatory tangent space projectors (16) by considering multi-indices that satisfy the nested conditions
@@
To prove that (30) is sufficient for the oblique tangent space projector (16) to have a cross interpolation property we have the following Lemma.
@@
$\begin{split}&\sum_{k=1}^{d}\left[P_{\leq k-1}P_{>k}Z\right]\left(\mathcal{I}^{\leq j-1},i_{j},\mathcal{I}^{>j}\right)=\\ &\qquad\left(\sum_{k=1}^{j-1}\left[P_{\leq k}P_{>k}Z\right]\left(\mathcal{I}^{\leq j-1},:,\mathcal{I}^{>j}\right)\right)+Z\left(\mathcal{I}^{\leq j-1},i_{j},\mathcal{I}^{>j}\right)+\left(\sum_{l=j+1}^{d}\left[P_{\leq l-1}P_{>l-1}Z\right]\left(\mathcal{I}^{\leq j-1},i_{j},\mathcal{I}^{>j}\right)\right)\end{split}$ (37)
@@
The cross interpolation property (34) of the tangent space projector resembles the cross interpolation property of TT-cross approximation [47] with nested indices. Next, we describe a novel index selection algorithm for constructing well-conditioned interpolatory projectors onto the tangent space $T_{Y}\mathcal{M}_{\bm{r}}$. Then we show that the resulting indices parameterize $Y$ with tensor cross interpolation.
@@ 4 Index selection for oblique projectors and cross interpolation
In the preceding section we constructed tangent space projectors (16) with the cross interpolation property (34) from oblique projectors (14) onto the bases $\bm{U}_{\leq k}$ and $\bm{V}_{>k}$. We now devise an efficient algorithm based on the discrete empirical interpolation method (DEIM) for computing indices $\{I^{\leq k},I^{>k}\}$, or equivalently multi-indices $\{\mathcal{I}^{\leq k},\mathcal{I}^{>k}\}$, that yield well-conditioned oblique projectors (14) defined by the matrices (27). The DEIM (recalled in Algorithm 1) is a greedy algorithm for computing indices that yield well-conditioned interpolatory projectors onto a given basis, e.g., $\bm{U}_{\leq k}$ or $\bm{V}_{>k}$. However we can not apply such algorithm directly to the matrices $\bm{U}_{\leq k},\bm{V}_{>k}$ as they have dimensions $(n_{1}\cdots n_{k})\times r_{k}$ and $(n_{k+1}\cdots n_{d})\times r_{k}$, respectively, which is too large to store in memory and process with the DEIM algorithm. To address the problem of memory, we do not store the matrices $\bm{U}_{\leq k}$ and $\bm{V}_{>k}$ directly but rather the left orthogonal cores $U_{j}$ and right orthogonal cores $V_{j}$ that can be used to construct these matrices
@@
To address the computational cost we propose an algorithm that only samples from a small subset of entries of the matrices $\bm{U}_{\leq k},\bm{V}_{>k}$. The key idea is to compute the indices $I^{\leq k}$ recursively for $k=1,2,\ldots,d-1$ by sampling from $\bm{U}_{\leq k}$ only indices corresponding to multi-indices $\mathcal{I}^{\leq k}$ that are nested (30). By considering only nested indices we reduce the number of possible indices $i_{1}\cdots i_{k}\in I^{\leq k}$ from $n_{1}\cdots n_{k}$ to $r_{k-1}n_{k}$. We use the same idea to construct the nested index sets $\mathcal{I}^{>k}$ sequentially for $k=d-1,d-2,\ldots,1$. Remarkably, it is always possible to obtain well-conditioned interpolatory projectors using this nested sampling approach. Moreover, since the indices obtained from this sampling approach are nested by construction, the resulting tangent space projector (16) has the cross interpolation property (34). Such cross interpolation property will be useful for the dynamical low-rank approximation schemes presented in Section 5. Hereafter we present the nested index selection algorithm in detail and show that the resulting indices yield well-conditioned interpolatory projectors.
Algorithm 1 DEIM index selection (adapted from [49])
1:$\bm{V}\in\mathbb{R}^{n\times r}$ with $n\geq r$
2:$\bm{l}$, a vector with $r$ distinct indices from $\{1,\ldots,n\}$
3:$\bm{v}=\bm{V}(:,1)$
[…]
5:for $j=2,3,\ldots,r$  do
6:     $\bm{v}=\bm{V}(:,j)$
7:     $\bm{c}=\bm{V}(l,1:j-1)^{-1}\bm{v}(\bm{l})$
8:     $\bm{r}=\bm{v}-\bm{V}(:,1:j-1)\bm{c}$
9:     $[~{},l_{j}]=\max(|\bm{r}|)$
10:     $\bm{l}=[\bm{l};l_{j}]$
11:end for
@@ 4.1 The TT-cross-DEIM algorithm
Let us denote by $\texttt{DEIM}(\cdot)$ a subroutine that takes a matrix of size $n\times r$ as input and outputs a collection of $r$ indices computed with the DEIM index selection Algorithm 1. To compute the indices $I^{\leq j}$, begin by applying the DEIM algorithm to the $n_{1}\times r_{1}$ matrix $\bm{U}_{\leq 1}$
@@
which corresponds to $I^{\leq j}$. This procedure computes index sets $I^{\leq j}$ sequentially for $j=1,2,\ldots,d-1$ by applying the DEIM sampling algorithm to the restricted matrices $\widehat{\bm{U}}_{\leq j}$ of dimension $r_{j-1}n_{j}\times r_{j}$. In practice the restricted matrices $\widehat{\bm{U}}_{\leq j}$ are not obtained from $\bm{U}_{\leq j}$ as written in (44) but rather from the low-dimensional tensor cores $U_{j}$ that construct $\bm{U}_{\leq j}$ in (39). The computation of $I^{\leq j}$ from the tensor cores $U_{j}$ is depicted in Figure 3 with $d=4$.
@@
The entire algorithm is summarized in Algorithm 2 and a tensor network diagram of the left-to-right sweep for computing $I^{\leq j}$ from the TT-cores $U_{j}$ is shown in Figure 3. In Algorithm 2 we denote by ind2sub the Matlab function that reshapes linear indices to multi-indices. As mentioned above, the multi-index sets ${\mathcal{I}^{\leq k},\mathcal{I}^{>k}}$ obtained in (46) and (50) are nested (30) by construction. Thus the oblique tangent space projector (16) constructed from these indices is a cross interpolant (see Theorem 3.1), provided it is well-defined. It will be shown shortly that such projector is in fact well-defined.
@@ 4.2 Condition of the oblique projectors
The next result shows that the TT-cross-DEIM produces indices that yield invertible matrices (51) and therefore define oblique projectors (14) and (16).
@@
proving the result for $\bm{M}_{j}$. The statement for $\bm{N}_{j}$ is proven similary with induction on $j=d-1,d-2,\ldots,1$. $\square$
In addition to producing invertible matrices (51), the TT-cross-DEIM is a greedy algorithm that keeps the condition number of such matrices as small as possible subject to the constraint that the indices are nested. Indeed, the DEIM algorithm is employed to obtain $\bm{l}_{\leq j}$ in (45) and $\bm{l}_{>j}$ in (49) which greedily minimizes the condition number of $\widehat{\bm{U}}_{\leq j}(\bm{l}_{\leq j},:)$ and $\widehat{\bm{V}}_{>j}\left(\bm{l}_{>j},:\right)$. Given that $\widehat{\bm{U}}_{\leq j}$ is the restriction of $\bm{U}_{\leq j}$ to indices $I^{\leq j-1}$ (see (44)) and $\widehat{\bm{V}}_{>j}$ is the restriction of $\bm{V}_{>j}$ to indices $I^{>j+1}$ (see (48)), the TT-cross-DEIM does in fact select $\{I^{\leq j},I^{>j}\}$ to minimize the condition number of (51) subject to the constraint that the corresponding multi-indices are nested (30).
@@ 4.3 Tensor cross interpolation
We have shown above that the multi-indices obtained with the TT-cross-DEIM yield a well-conditioned interpolatory projector (16) onto the tangent space $T_{Y}\mathcal{M}_{\bm{r}}$. Incidentally, we can use the same multi-index sets to parameterize $Y$ with tensor cross interpolation. Recall that TT-cross approximation [14, 37, 47, 40] is a specific instance of the TT format (3) that generalizes matrix CUR decomposition to tensors. In this representation the TT-cores are defined by the entries of $Y$
@@
Several algorithms for computing tensor cross approximations (54) from black-box tensors based on the maximum volume principle have recently been developed [37, 14]. The purpose of our proposed TT-cross-DEIM algorithm is to obtain well-conditioned interpolatory tangent space projections for a tensor $Y\in\mathcal{M}_{\bm{r}}$, not for black-box tensor approximation. However, it was recently demonstrated in [20] that DEIM-based cross approximation algorithms can be applied iteratively to obtain tensor cross approximations from black-box tensors with comparable performance to the corresponding maximum volume algorithms.
5 Time integration on low-rank tensor manifolds
@@
In this section we introduce dynamical low-rank schemes based on oblique tangent space projections for approximating the nonlinear tensor differential equation (2) on the TT manifold (8). The idea of dynamical low-rank methods is to integrate an approximate solution $Y(t)$ on $\mathcal{M}_{\bm{r}}$ by projecting $G(Y,t)$ defining the ODE (2) onto the tangent space $T_{Y}\mathcal{M}_{\bm{r}}$ at each time $t$. Conventionally, dynamical low-rank methods have employed orthogonal projectors (13) to obtain the best approximation of $G(Y,t)$ in the tangent space (see Figure 2). However, as we mentioned previously, the computational cost of orthogonal projection onto the tangent space can be $\mathcal{O}(n^{d})$ when $G$ lacks low-rank structure. We now propose new dynamical low-rank methods based on oblique tangent space projections that have cost scaling as $\mathcal{O}(dnr^{3})$ in for many applications where methods based on orthogonal projections cost $\mathcal{O}(n^{d})$. The proposed oblique dynamical low-rank methods are efficient whenever it is possible to evaluate $G(X,t)$ entry-wise and do not require $G$ to have any low-rank structure.
If the initial condition $X_{0}$ in (2) belongs to $\mathcal{M}_{\bm{r}}$, then the solution to the projected tensor differential equation
$\frac{dY(t)}{dt}=P_{Y(t)}G(Y(t),t),\qquad Y(0)=X_{0},$ (57)
@@
is a dynamical low-rank approximation of the solution to (2) that remains on $\mathcal{M}_{\bm{r}}$ for all $t$. Hereafter we propose two time integration schemes for solving the dynamical low-rank equation (57) with interpolatory tangent space projectors (16). The first scheme, referred to as TT-cross time integration, integrates forward in time the entries of the solution tensor $Y(t)$ required to construct the TT-cross interpolant (54) at any time $t$. The second scheme is a direct generalization of the projector-splitting integrator for orthogonal dynamical low-rank approximation introduced in [31] for matrices and subsequently generalized to TTs [32], Tucker tensors [30] and tree tensor networks [7], to oblique tangent space projectors in the TT format.
5.1 TT-cross time integration

### v3

@@ 3.1 Orthogonalization of tensor trains
Orthogonal TT representations are fundamental for executing many operations in the TT format. We will use them in this paper to obtain projectors onto the tangent spaces of TT manifolds. Hereafter, we recall an algorithm for orthogonalizing TTs by recursively applying QR-decomposition to TT-core unfoldings. Begin by taking a QR-decomposition of the left unfolding of $C_{1}$
@@
$Y=U_{\leq k}\bm{S}_{k}V_{>k}.$ (15)
Utilizing the unfolding matrices of partial products (14) we also have a decomposition of the $k$th unfolding matrix
@@ 4 Projections onto the TT tangent space
is a smooth embedded submanifold of $\mathbb{R}^{n_{1}\times\cdots\times n_{d}}$ [27]. Hence for any tensor $Y\in\mathcal{M}_{\bm{r}}$ we can define the tangent space $T_{Y}\mathcal{M}_{\bm{r}}$, which is a vector subspace of $\mathbb{R}^{n_{1}\times\cdots\times n_{d}}$ that linearizes the manifold $\mathcal{M}_{\bm{r}}$ around $Y$. Given a TT representation (12) of $Y$, any element of the tangent space can be written (non-uniquely) as
@@ 4.1 The orthogonal tangent space projector
onto the range and co-range of $\bm{Y}^{\langle k\rangle}$. The orthogonal bases $\bm{U}_{\leq k}$ and $\bm{V}_{>k}$ for the range and co-range of $\bm{Y}^{\langle k\rangle}$ can be obtained from the TT-orthogonalization procedure described in Section 3.1. These projectors act on the matrix space $\mathbb{R}^{(n_{1}\cdots n_{k})\times(n_{k+1}\cdots n_{d})}$. To construct the orthogonal tangent space projector $\widehat{P}_{Y}$ it is convenient to define projectors corresponding to (19) that act on the tensor space $\mathbb{R}^{n_{1}\times\cdots\times n_{d}}$ by
@@ 4.2 Oblique tangent space projectors
with $P_{\leq 0}=1$, defined for any tensor $Z\in\mathbb{R}^{n_{1}\times\cdots\times n_{d}}$, is an oblique projector onto the tangent space $T_{Y}\mathcal{M}_{\bm{r}}$.
@@
which we write in the TT format as
@@
We remark that the orthogonality of the bases $\bm{U}_{\leq k}$ and $\bm{V}_{>k}$ is not necessary for the construction of oblique tangent space projectors (24). However, imposing such orthogonality is advantageous for the numerical stability and accuracy of the projectors.
@@ 4.3 Interpolatory tangent space projectors
Next we introduce a special class of oblique tangent space projectors with a cross interpolation property that enables efficient dynamical low-rank approximation. Such oblique projectors are obtained by selecting $\bm{X}_{\leq k}$ and $\bm{X}_{>k}$ in (24) as $r_{k}$ columns of the identity matrix with appropriate dimension
@@
In order to obtain an oblique tangent space projectors (24) that interpolate, we consider multi-indices that satisfy the nested conditions
@@
To prove that (38) is sufficient for the oblique tangent space projector (24) to interpolate, we have the following Lemma.
@@
$\begin{split}&\sum_{k=1}^{d}\left[P_{\leq k-1}P_{>k}Z\right]\left(\mathcal{I}^{\leq j-1},i_{j},\mathcal{I}^{>j}\right)=\\ &\qquad\left(\sum_{k=1}^{j-1}\left[P_{\leq k}P_{>k}Z\right]\left(\mathcal{I}^{\leq j-1},:,\mathcal{I}^{>j}\right)\right)+Z\left(\mathcal{I}^{\leq j-1},i_{j},\mathcal{I}^{>j}\right)+\left(\sum_{l=j+1}^{d}\left[P_{\leq l-1}P_{>l-1}Z\right]\left(\mathcal{I}^{\leq j-1},i_{j},\mathcal{I}^{>j}\right)\right).\end{split}$ (45)
@@
The cross interpolation property (42) of the tangent space projector resembles the interpolation property of TT-cross approximation [47] with nested indices. Next, we describe a novel index selection algorithm for constructing interpolatory projectors onto the tangent space $T_{Y}\mathcal{M}_{\bm{r}}$.
@@ 5 Index selection for oblique projectors and cross interpolation
In the preceding section we constructed tangent space projectors (24) with the cross interpolation property (42) from oblique projectors (22) onto the bases $\bm{U}_{\leq k}$ and $\bm{V}_{>k}$. We now devise an efficient algorithm based on the discrete empirical interpolation method (DEIM) for computing indices $\{I^{\leq k},I^{>k}\}$, or equivalently multi-indices $\{\mathcal{I}^{\leq k},\mathcal{I}^{>k}\}$, that yield well-defined interpolatory projectors (22) defined by the matrices (35). The DEIM (recalled in Algorithm 1) greedily selects indices to minimize the condition number of interpolatory projectors onto a given basis, e.g., $\bm{U}_{\leq k}$ or $\bm{V}_{>k}$, as much as possible. However we can not apply such algorithm directly to the matrices $\bm{U}_{\leq k},\bm{V}_{>k}$ as they have dimensions $(n_{1}\cdots n_{k})\times r_{k}$ and $(n_{k+1}\cdots n_{d})\times r_{k}$, respectively, which is too large to store in memory and process with the DEIM algorithm. To address the problem of memory, we do not store the matrices $\bm{U}_{\leq k}$ and $\bm{V}_{>k}$ directly but rather the left orthogonal cores $U_{j}$ and right orthogonal cores $V_{j}$ that can be used to construct these matrices
@@
To address the computational cost we propose an algorithm that only samples from a small subset of entries of the matrices $\bm{U}_{\leq k},\bm{V}_{>k}$. The key idea is to compute the indices $I^{\leq k}$ recursively for $k=1,2,\ldots,d-1$ by sampling from $\bm{U}_{\leq k}$ only indices corresponding to multi-indices $\mathcal{I}^{\leq k}$ that are nested (38). By considering only nested indices we reduce the number of possible indices $i_{1}\cdots i_{k}\in I^{\leq k}$ from $n_{1}\cdots n_{k}$ to $r_{k-1}n_{k}$. We use the same idea to construct the nested index sets $\mathcal{I}^{>k}$ sequentially for $k=d-1,d-2,\ldots,1$. Since the indices obtained from this sampling approach are nested by construction, the resulting tangent space projector (24) has the cross interpolation property (42). Such cross interpolation property will be useful for the dynamical low-rank approximation schemes presented in Section 6. Hereafter we present the nested index selection algorithm in detail and then show in Theorem 5.1 that this nested sampling method always produces well-defined interpolatory projectors.
@@ 5.1 The TT-cross-DEIM algorithm
To compute the indices $I^{\leq j}$, begin by applying the DEIM algorithm to the $n_{1}\times r_{1}$ matrix $\bm{U}_{\leq 1}$
@@
which corresponds to $I^{\leq j}$. This procedure computes index sets $I^{\leq j}$ sequentially for $j=1,2,\ldots,d-1$ by applying the DEIM sampling algorithm to the restricted matrices $\widehat{\bm{U}}_{\leq j}$ of dimension $r_{j-1}n_{j}\times r_{j}$. In practice the restricted matrices $\widehat{\bm{U}}_{\leq j}$ are not obtained from $\bm{U}_{\leq j}$ as written in (52) but rather from the low-dimensional tensor cores $U_{j}$ that construct $\bm{U}_{\leq j}$ in (47).
@@
The entire algorithm is summarized in Algorithm 2 and a tensor network diagram of the left-to-right sweep for computing $I^{\leq j}$ from the TT-cores $U_{j}$ is shown in Figure 3 with $d=4$. In Algorithm 2 we denote by ind2sub the Matlab function that reshapes linear indices to multi-indices. As mentioned above, the multi-index sets ${\mathcal{I}^{\leq k},\mathcal{I}^{>k}}$ obtained in (54) and (58) are nested (38) by construction. Thus the oblique tangent space projector (24) constructed from these indices is a cross interpolant (see Theorem 4.1), provided it is well-defined. It is shown in Section 5.2 that such projector is in fact well-defined.
@@ 5.2 Condition of the oblique projectors
The following result shows that the TT-cross-DEIM produces indices that yield invertible matrices (59) and therefore define oblique projectors (22) and (24).
@@
proving the result for $\bm{M}_{j}$. The statement for $\bm{N}_{j}$ is proven similarly with induction on $j=d-1,d-2,\ldots,1$. $\square$
In addition to generating invertible matrices (59), the TT-cross-DEIM is a greedy algorithm that aims to minimize the condition number of these matrices as much as possible while ensuring the indices remain nested. Indeed, the DEIM algorithm is used to compute $\bm{l}_{\leq j}$ in (53) and $\bm{l}_{>j}$ in (57), selecting indices greedily to keep the condition numbers of $\widehat{\bm{U}}_{\leq j}(\bm{l}{\leq j},:)$ and $\widehat{\bm{V}}_{>j}(\bm{l}{>j},:)$ small. Moreover, $\widehat{\bm{U}}_{\leq j}$ is the restriction of $\bm{U}_{\leq j}$ to the indices $I^{\leq j-1}$ (see (52)), and $\widehat{\bm{V}}_{>j}$ is the restriction of $\bm{V}_{>j}$ to the indices $I^{>j+1}$ (see (56)). We also note that other sparse sampling methods can be used in place of DEIM in the TT-cross-DEIM algorithm, e.g., Q-DEIM or oversampling methods, which can yield better conditioned interpolatory projectors in some cases.
@@ 5.3 Tensor cross interpolation
We have shown above that the multi-indices obtained with the TT-cross-DEIM produce a well-defined interpolatory projector (24) onto the tangent space. Incidentally, we can use the same multi-index sets to parameterize $Y$ with tensor cross interpolation. Recall that TT-cross approximation [14, 37, 47, 40] is a specific instance of the TT format (12) that generalizes the matrix CUR decomposition to tensors. In this representation the TT-cores are defined by the entries of $Y$
@@
Several algorithms for computing tensor cross approximations (62) from black-box tensors based on the maximum volume principle have recently been developed [37, 14]. The purpose of our proposed TT-cross-DEIM algorithm is to obtain interpolatory tangent space projections for a tensor $Y\in\mathcal{M}_{\bm{r}}$, not for black-box tensor approximation. However, it was recently demonstrated in [20] that DEIM-based cross approximation algorithms can be applied iteratively to obtain tensor cross approximations from black-box tensors with comparable performance to the corresponding maximum volume algorithms.
6 Time integration on tensor train manifolds
@@
We now consider the dynamical low-rank evolution equation (3) for tensors ($d\geq 2$) using interpolatory projectors (24) onto tangent spaces of TT manifolds. The concept of dynamical low-rank tensor approximation is a natural extension the dynamical low-rank matrix approximation described in Section 2. Similar to the matrix case, classical dynamical low-rank tensor approximation uses the orthogonal projector (21) to obtain the best approximation (in the Frobenius norm) of $G(Y,t)$ in the tangent space of the TT manifold (see Figure 1). However, as noted earlier, orthogonal projection onto the TT tangent space can have a computational cost $\mathcal{O}(n^{d})$ when $G$ lacks low-rank structure. By replacing the orthogonal projector with an interpolatory projector onto the TT tangent space we propose new dynamical low-rank methods with computational cost scaling as $\mathcal{O}(dnr^{3})$ for a large class of nonlinear functions $G$ that do not have rank structure. In particular, the proposed interpolatory dynamical low-rank tensor methods are efficient whenever it is possible to evaluate the tensor $G(X,t)$ entry-wise.
We also point out that cross approximation algorithms based on the maximum volume principle developed in [37] are designed to obtain TT approximations of tensors that can be evaluated entry-wise. TT-cross based on maximum volume can be used to obtain a low-rank approximation of the tensor $G(Y,t)$ at each time step. Such approximation can then be projected orthogonally onto the tangent space for a dynamical low-rank method or used in a step-truncation scheme. The TT-cross-DEIM index selection strategy developed in the present work differs in that it selects interpolation indices from the solution tensor $Y(t)$, not from $G(Y,t)$. Such indices are selected so that $Y$ can be represented using a TT-cross interpolant. More importantly, it allows $G(Y,t)$ to be interpolated directly onto the tangent space of the TT manifold at $Y$, making the TT-cross-DEIM particularly suitable for efficient dynamical low-rank approximation.
@@
Hereafter we propose two time integration schemes for solving the dynamical low-rank equation (3) with interpolatory TT tangent space projectors (24). The first scheme, referred to as TT-cross time integration, extends the matrix cross integrator described in Section 2.1.1 to tensors in the TT format. This method integrates forward in time the entries of the solution tensor $Y(t)$ required to construct the TT-cross interpolant (62) at any time $t$. The second scheme extends the projector-splitting scheme described in Section 2.1.2 to tensors in the TT format. It is a direct generalization of the projector-splitting integrator for orthogonal dynamical low-rank approximation introduced in [31] for matrices and subsequently generalized to TTs [32], Tucker tensors [30] and tree tensor networks [7], to interpolatory tangent space projectors in the TT format.
6.1 TT-cross integrator